OpenStudy (anonymous):

Why does the mass spectrum of Br2 contain three signals whose heights are almost in the ratio of 1:2:1? What are the origins of these signals? Why does the spectrum contain two other signals of roughly equal height? What are the origins of these signals?

OpenStudy (anonymous):

If we assume the proportion of the two bromine isotopes 79-Br and 81-Br are each 50% then the combinations to diatomic molecules are as follows: (79-Br)-(79Br) [combined mass = 158, contribution 25%] (79-Br)-(81-Br) [combined mass = 160, contribution 25%] (81-Br)-(79-Br) [combined mass = 160, contribution 25%] (81-Br)-(81Br) [combined mass = 162, contribution 25%] You will note that two combinations have the same mass, giving the 1:2:1 ratio.

OpenStudy (abb0t):

Those signals are of isotopes of Br\(_2\).